### v1

@@ 1 Introduction
Therefore, to overcome the respective limitations of flow matching and BFNs, we integrate their core concepts and introduce a novel and versatile parameter interpolation flow (PIF) model, as illustrated in Fig.  1. Our framework treats the complete data distribution as a superposition of Dirac distributions, each corresponding to an individual data point, and trains the model to learn the underlying distributional parameters. Specifically, PIF employs a probabilistic framework that interpolates between Dirac distributions and a chosen prior (e.g., Gaussian or Dirichlet distributions), thereby defining a continuous trajectory toward the prior. During training, the model learns to predict the parameters of intermediate distributions along this trajectory, conditioned on a randomly sampled timestep $t$, and is optimized via Kullback-Leibler (KL) divergence. At inference time, the process begins from the prior and iteratively refines predictions over predefined steps, gradually converging toward the target data distribution. Compared with conventional generative models, PIF facilitates smooth transformations in the parameter space of distributions, rendering it highly versatile for both continuous and discrete data domains. Furthermore, PIF offers flexible prior selection, enabling adaptation to diverse tasks without requiring complex closed-form derivations.
@@ 2.4.2 Analysis of MolPIF on local geometries
Figure 2: Local geometry analysis. a, Ratio of different-sized rings generated by the model. b, c, d, Jensen-Shannon divergence (JSD) of bond lengths, bond angles and torsion angles distributions between generated molecules and the reference molecules in the test set.
@@
In addition to conventional properties, we evaluated the local geometric characteristics of the generated molecules through (1) statistical comparison of 3- to 8-membered ring frequencies between reference and generated molecules, and (2) substructural level’s Jensen-Shannon divergence (JSD) analysis of key geometric parameters including common covalent bond lengths (CC, C=C, CO, CN, C=N, OP, C=O), bond angles (CCC, C=C=C, CCO, C=C=N, CCN), and torsion angles (CCCC, C=C=C=C, CCOC, CCCO) distributions (Fig.  2). This comprehensive analysis provided insights into the model’s ability to reproduce the precise geometric features of molecular structures.
@@
Fig. 2a presents the occurrence frequencies of 3- to 8-membered rings in both model-generated molecules and reference molecules. MolPIF demonstrated significantly lower proportions of unstable small rings (3- and 4-membered) in its generated molecules. Notably, no 3-membered rings were observed in MolPIF’s outputs, while 4-membered rings occurred at only 0.44% frequency, which was substantially lower than diffusion-model-based baselines. MolPIF exhibited a strong preference for generating 5- and 6-membered rings, which are widely employed in drug design. Specifically, 6-membered rings accounted for 76.88% of generated structures. For larger 7- and 8-membered rings, MolPIF maintained relatively low generation frequencies: 7-membered ring production was comparable to MolCRAFT and slightly higher than AR and Pocket2Mol, while 8-membered rings showed the lowest occurrence among all compared methods (Extended Data Table 3). The AR, constrained by its local topology generation mechanism, tended to overproduce small ring structures. Diffusion-based models TargetDiff and DecompDiff displayed relatively higher frequencies for 4-, 7-, and 8-membered rings but underperformed in generating the most prevalent 6-membered rings, indicating their difficulty in fitting the reference molecular ring distribution during training. In contrast, both MolCRAFT and MolPIF, which operate in parameter space for molecular generation, produced ring distributions that closely approximated the reference molecular profiles.
Fig. 2b-d presents the JSD of bond lengths, bond angles, and torsion angles between molecules generated by MolPIF and baseline methods compared to reference molecules in the test set. Overall, the JSD between covalent bond lengths, bond angles, and torsion angles of model-generated molecules and the reference set followed a consistent trend: autoregressive models >diffusion-based models >parametric generative models. Among parametric generative models, MolPIF demonstrated lower JSD than MolCRAFT for bond lengths of C=C and C=N bonds. The unmasked variant of MolPIF (MolPIF w/o mask) outperformed MolCRAFT in bond length JSD for all bonds except CN and OP. For bond angle distributions, MolPIF showed slightly inferior performance to MolCRAFT, though both MolPIF and its unmasked variant maintained relatively low JSD values overall. In torsion angle distributions, no significant differences were observed except for C=C=C=C—where MolCRAFT performed best—while MolPIF (w/o mask) achieved optimal JSD values for all other torsion angles analyzed. Detailed substructure-level metrics are provided in Extended Data Table 456.
@@ 2.4.4 Analysis of the prior distribution selection
Regarding the performance of the model in generating molecular substructures, Fig. 4 presents a visualization of the specific distributions for selected representative cases, with detailed data available in Extended Data Tables 78, and 9. The results were compared with MolCRAFT, which demonstrated exceptional performance in molecular substructure generation. For bond lengths, MolPIF showed superior capability in capturing the peak values of reference molecular bond length distributions compared to MolCRAFT. The fitted distributions by MolPIF aligned more closely with the reference bond length distributions (Fig. 4a-c), exhibiting generally lower JSD values (Extended Data Table 7). Notably, the Laplace prior distribution provided a more pronounced advantage, particularly for MolPIF(La), which produced distributions that most closely matched the original among all models. For C=O bonds, MolPIF(La) successfully fit three distinct modes that other models failed to capture (Fig. 4c). In terms of bond angles, MolPIF(La) also demonstrated outstanding performance, achieving the lowest JSD values for C=C=C and CCO distributions compared to the reference molecules, while attaining the second best JSD performance for CCC and CCN (Extended Data Table 8). Taking CCO as an example, MolPIF(La)-generated molecules better approximated the true distribution, with more accurate peak positions and more obvious curve variations than other models (Fig. 4d). Regarding torsion angles, MolPIF(w/o mask) exhibited superior performance, achieving optimal JSD values for CCCC, CCOC, and CCCO, while MolPIF(La) attained the second best performance for C=C=C=C and CCOC (Extended Data Table 9). However, none of the models performed exceptionally well in fitting specific torsion angle modes. For instance, in the case of CCCC, all models only partially captured the distribution modes, with some deviation in peak positions (Fig. 4e).
@@ 2.4.5 The performance of MolPIF in lead optimization
Across all subtasks involving 1umd, 3ZCW, and 6KZZ, a substantial proportion of the MolPIF-generated molecules demonstrated superior performance compared to the reference molecules in multiple key metrics (Extended Data Tables 10,  11). These findings indicated that MolPIF not only enabled effective structural optimization but also maintained or enhanced molecules’ 3D spatial compatibility and binding capability. Under identical experimental conditions, MolPIF generated a higher proportion of molecules with properties superior to reference compounds compared to Delete and DeepFrag, underscoring its broad applicability and substantial potential in lead optimization tasks.
@@ Appendix B Extended Data
Table 3: Ratio of different-sized rings generated by models
3 4 5 6 7 8
AR 0.3086 0.0030 0.1556 0.4926 0.0190 0.0082
Pocket2Mol 0.0012 0.0002 0.1626 0.7983 0.0259 0.0034
TargetDiff 0.0000 0.0270 0.2971 0.4896 0.1170 0.0259
DecompDiff 0.0264 0.0391 0.3425 0.4396 0.1135 0.0178
MolCRAFT 0.0000 0.0022 0.2310 0.6986 0.0540 0.0062
MolPIF 0.0000 0.0044 0.1597 0.7688 0.0565 0.0034
MolPIF(w/o mask) 0.0000 0.0045 0.1918 0.7269 0.0614 0.0059
Test set 0.0172 0.0000 0.2961 0.6609 0.0086 0.0000
Table 4: JSD of bond lengths between reference and generated molecules
CC C=C CO CN C=N OP C=O
AR 0.6098 0.4503 0.4904 0.4722 0.5514 0.6762 0.5563
Pocket2Mol 0.4937 0.4144 0.4524 0.4222 0.4840 0.5257 0.5095
TargetDiff 0.3690 0.2634 0.4208 0.3623 0.2349 0.4430 0.4608
DecompDiff 0.3265 0.2378 0.3467 0.3374 0.2344 0.4524 0.3903
MolCRAFT 0.3206 0.3231 0.3536 0.3010 0.2497 0.3410 0.3339
MolPIF 0.3865 0.1978 0.4104 0.3425 0.2068 0.3659 0.3351
MolPIF(w/o mask) 0.2885 0.2257 0.3475 0.3161 0.2216 0.3505 0.3278
Table 5: JSD of bond angles between reference and generated molecules
CCC C=C=C CCO C=C=N CCN
AR 0.3721 0.5719 0.4770 0.5366 0.4474
Pocket2Mol 0.3799 0.4804 0.4753 0.5060 0.4432
TargetDiff 0.3450 0.2828 0.4396 0.4537 0.4374
DecompDiff 0.3383 0.2549 0.3898 0.4288 0.4179
MolCRAFT 0.3015 0.1741 0.3473 0.4508 0.3796
MolPIF 0.3716 0.2078 0.4349 0.4323 0.4178
MolPIF(w/o mask) 0.3073 0.2219 0.3751 0.4509 0.4099
Table 6: JSD of torsion angles between reference and generated molecules
CCCC C=C=C=C CCOC CCCO
AR 0.3778 0.7130 0.3736 0.4270
Pocket2Mol 0.3220 0.5339 0.3683 0.4009
TargetDiff 0.3134 0.3730 0.3416 0.4030
DecompDiff 0.3455 0.2403 0.3423 0.4113
MolCRAFT 0.2818 0.1555 0.3417 0.3977
MolPIF 0.2901 0.3022 0.3417 0.3911
MolPIF(w/o mask) 0.2676 0.2667 0.3337 0.3853

### v2

@@ 1 Introduction
Therefore, to overcome the respective limitations of flow matching and BFNs, we introduce a novel and versatile parameter interpolation flow (PIF) model, as illustrated in Fig.  1. Our framework treats the complete data distribution as a superposition of Dirac distributions, each corresponding to an individual data point, and trains the model to learn the underlying distributional parameters. Specifically, PIF employs a probabilistic framework that interpolates between Dirac distributions and a chosen prior (e.g., Gaussian or Dirichlet distributions), thereby defining a continuous trajectory toward the prior. During training, the model learns to predict the parameters of intermediate distributions along this trajectory, conditioned on a randomly sampled timestep $t$, and is optimized via Kullback-Leibler (KL) divergence. At inference time, the process begins from the prior and iteratively refines predictions over predefined steps, gradually converging toward the target data distribution. Compared with conventional generative models, PIF facilitates smooth transformations in the parameter space of distributions, rendering it highly versatile for both continuous and discrete data domains. Furthermore, PIF offers flexible prior selection, enabling adaptation to diverse tasks without requiring complex closed-form derivations.
@@ 2.4.2 Analysis of MolPIF on local geometries
In addition to conventional properties, we evaluated the local geometric characteristics of the generated molecules through (1) statistical comparison of 3- to 8-membered ring frequencies between reference and generated molecules (Table 2) , and (2) substructural level’s Jensen-Shannon divergence (JSD) analysis of key geometric parameters including common covalent bond lengths (CC, C=C, CO, CN, C=N, OP, C=O), bond angles (CCC, C=C=C, CCO, C=C=N, CCN), and torsion angles (CCCC, C=C=C=C, CCOC, CCCO) distributions. This comprehensive analysis provided insights into the model’s ability to reproduce the precise geometric features of molecular structures.
@@
Table 2 presents the occurrence frequencies of 3- to 8-membered rings in both model-generated molecules and reference molecules. MolPIF demonstrated significantly lower proportions of unstable small rings (3- and 4-membered) in its generated molecules. Notably, no 3-membered rings were observed in MolPIF’s outputs, while 4-membered rings occurred at only 0.44% frequency, which was substantially lower than diffusion-model-based baselines. MolPIF exhibited a strong preference for generating 5- and 6-membered rings, which are widely employed in drug design. Specifically, 6-membered rings accounted for 76.88% of generated structures. For larger 7- and 8-membered rings, MolPIF maintained relatively low generation frequencies: 7-membered ring production was comparable to MolCRAFT and slightly higher than AR and Pocket2Mol, while 8-membered rings showed the lowest occurrence among all compared methods. The AR, constrained by its local topology generation mechanism, tended to overproduce small ring structures. Diffusion-based models TargetDiff and DecompDiff displayed relatively higher frequencies for 4-, 7-, and 8-membered rings but underperformed in generating the most prevalent 6-membered rings, indicating their difficulty in fitting the reference molecular ring distribution during training. In contrast, both MolCRAFT and MolPIF, which operate in parameter space for molecular generation, produced ring distributions that closely approximated the reference molecular profiles.
For the JSD of bond lengths, bond angles, and torsion angles between molecules generated by MolPIF and baseline methods compared to reference molecules in the test set. Overall, the JSD between covalent bond lengths, bond angles, and torsion angles of model-generated molecules and the reference set followed a consistent trend: autoregressive models >diffusion-based models >parametric generative models, which aligned with the conclusions demonstrated in MolCRAFT [29]. The specific performance of the superior-performing parametric generative models regarding the JSD of bond lengths, bond angles, and torsion angles will be discussed in detail in Section 2.4.4.
@@
Table 2: Ratio of different-sized rings generated by models
3 4 5 6 7 8
AR 0.3086 0.0030 0.1556 0.4926 0.0190 0.0082
Pocket2Mol 0.0012 0.0002 0.1626 0.7983 0.0259 0.0034
TargetDiff 0.0000 0.0270 0.2971 0.4896 0.1170 0.0259
DecompDiff 0.0264 0.0391 0.3425 0.4396 0.1135 0.0178
MolCRAFT 0.0000 0.0022 0.2310 0.6986 0.0540 0.0062
MolPIF 0.0000 0.0044 0.1597 0.7688 0.0565 0.0034
MolPIF(w/o mask) 0.0000 0.0045 0.1918 0.7269 0.0614 0.0059
Test set 0.0172 0.0000 0.2961 0.6609 0.0086 0.0000
@@ 2.4.4 Analysis of the prior distribution selection
Regarding the performance of the model in generating molecular substructures, Fig. 3 presents a visualization of the specific distributions for selected representative cases, with detailed data available in Extended Data Tables 45, and 6. The results were compared with MolCRAFT, which demonstrated exceptional performance in molecular substructure generation. For bond lengths, MolPIF showed superior capability in capturing the peak values of reference molecular bond length distributions compared to MolCRAFT. The fitted distributions by MolPIF aligned more closely with the reference bond length distributions (Fig. 3a-c), exhibiting generally lower JSD values (Extended Data Table 4). Notably, the Laplace prior distribution provided a more pronounced advantage, particularly for MolPIF(La), which produced distributions that most closely matched the original among all models. For C=O bonds, MolPIF(La) successfully fit three distinct modes that other models failed to capture (Fig. 3c). In terms of bond angles, MolPIF(La) also demonstrated outstanding performance, achieving the lowest JSD values for C=C=C and CCO distributions compared to the reference molecules, while attaining the second best JSD performance for CCC and CCN (Extended Data Table 5). Taking CCO as an example, MolPIF(La)-generated molecules better approximated the true distribution, with more accurate peak positions and more obvious curve variations than other models (Fig. 3d). Regarding torsion angles, MolPIF(w/o mask) exhibited superior performance, achieving optimal JSD values for CCCC, CCOC, and CCCO, while MolPIF(La) attained the second best performance for C=C=C=C and CCOC (Extended Data Table 6). However, none of the models performed exceptionally well in fitting specific torsion angle modes. For instance, in the case of CCCC, all models only partially captured the distribution modes, with some deviation in peak positions (Fig. 3e).
@@ 2.4.5 The performance of MolPIF in lead optimization
Across all subtasks involving 1umd, 3ZCW, and 6KZZ, a substantial proportion of the MolPIF-generated molecules demonstrated superior performance compared to the reference molecules in multiple key metrics (Extended Data Tables 7,  8). These findings indicated that MolPIF not only enabled effective structural optimization but also maintained or enhanced molecules’ 3D spatial compatibility and binding capability. Under identical experimental conditions, MolPIF generated a higher proportion of molecules with properties superior to reference compounds compared to Delete and DeepFrag, underscoring its broad applicability and substantial potential in lead optimization tasks.